Please answer........................................

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
5 is the only digit when multiplied by 3 gives 5 as unit digit
 
hence p = 5
 
and then q = 7
 
as 73 * 75 = 575
 
hence p + q = 5 + 7 = 12